OpenStudy (anonymous):

y/8 - 1/3 = y/12 + 5/6

OpenStudy (anonymous):

y=28

OpenStudy (anonymous):

how

Directrix (directrix):

y/8 - 1/3 = y/12 + 5/6 The denomators are 8, 3, 12, and 6. 24 is a multiple of each of them. Multiply 24 times each term to clear the fractions.

Directrix (directrix):

y/8 - 1/3 = y/12 + 5/6 ---> By 24 --> 3 y -8 = 2y + 20

Directrix (directrix):

Do you agree?

Directrix (directrix):

3 y -8 = 2y + 20 3y - 2y = 20 + 8 y = 28

Directrix (directrix):

Questions?